OpenStudy (anonymous):

My calculator died & I don't have a new one yet. Would someone please solve this for me. Thanks!! 1.3√ 160 = ____

OpenStudy (amistre64):

google it lol

OpenStudy (amistre64):

what factors of 160 are perfect squares?

OpenStudy (anonymous):

I tried to Google it no luck. Even tried to find an online root calculator... it didn't except the decimal place...wanted Whole number

OpenStudy (anonymous):

Oh come pls....its been almost 30 yrs since school

OpenStudy (amistre64):

so you simply need the whole number of this thing?

OpenStudy (amistre64):

it was 20 yrs for me lol

OpenStudy (anonymous):

Its part of a formula for calculating the DPN for a sailboat. For racing. its used for multiple class races to even out

OpenStudy (amistre64):

1.3 * sqrt(160 is what google will accept as the equation)

OpenStudy (anonymous):

it is 1.3 root of the Sail Area. which is 16o sq-ft

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

1.3 * sqrt(160) that is

OpenStudy (anonymous):

I enterd the sqrt symbol Thanks!!!

OpenStudy (anonymous):

1.3 * sqrt(160) = 16.4438438

OpenStudy (amistre64):

1.3 * sqrt(160) = 16.4438438 from the google :)

OpenStudy (anonymous):

so 1.3√160= 16.44

OpenStudy (anonymous):

Thanks for the help!!

OpenStudy (amistre64):

youre welcome:)

OpenStudy (anonymous):

Its 16.44..... All the Best :D

OpenStudy (anonymous):

OK......thanks! but I think that is wrong...... Will someone with a good old fashioned HP Reverse polish notaion calculator check this for me. I think it is actualy: 1.3rd root of 160 = 49.5994556 1.3 * sqrt(160) = 16.4438438 is really saying 1.3 times the square root of 160 which is like 12.65, so 12.65 x 1.3 = 16.44 OK so after 30 yrs I may not remember everything I learned in HS, but I do know enough to check alternate sources, and I do have common sense. Thanks for the help!

OpenStudy (anonymous):

OK I guess no one on here right now prob even knows what I'm talking about when I say an HP Reverse polish notaion eh? Well its simple a term we used in the 90's for the Hewlet Packard high end calclators where you input in rverse of standard calculators.

OpenStudy (amistre64):

you think its 1.3 as the root to the radical instead of a square root?

OpenStudy (anonymous):

yes! 1.3root of 160

OpenStudy (anonymous):

1.3√160

OpenStudy (anonymous):

not 1.3 x √160

OpenStudy (amistre64):

1' 3/10 = 13/10.. 160^(10/13) is the answer to that one then

OpenStudy (anonymous):

huh?

OpenStudy (amistre64):

160^(10 / 13) = 49.6

OpenStudy (amistre64):

radicals are fractional exponents

OpenStudy (anonymous):

yes 1.3rd root of 160 = 49.5994556

OpenStudy (amistre64):

\[\sqrt[a]{b} = b^{1/a}\]

OpenStudy (anonymous):

now you get the medal. Thanks! for confirming the first answers given were erronious. Have a great day!!

OpenStudy (amistre64):

\[\sqrt[5]{12^2} = 12^{2/5}\]

OpenStudy (amistre64):

:) youre welcome

OpenStudy (amistre64):

\[\sqrt[1.3]{160}=\sqrt[13/10]{160}=160^{\frac{1}{13/10}}=160^{10/13}\]

OpenStudy (anonymous):

but what is (DPN) if (DPN)=A+[B(DR)] DR=18.923 and A,B are Linear Regression Coefficients

OpenStudy (amistre64):

yipes!!.... what math you in that has that?

OpenStudy (anonymous):

or does the A & B require a data set to determine the answer....been to long since Statistics 103

OpenStudy (anonymous):

I'm 20 yrs past College

OpenStudy (anonymous):

that is Statistics

OpenStudy (amistre64):

ahh...im taking intro to stats this semester

OpenStudy (anonymous):

I could do it if I had my old calculator and the cheat sheet I think

OpenStudy (anonymous):

u in HS or College

OpenStudy (amistre64):

college at 40

OpenStudy (anonymous):

oh...LOL I went @ 27

OpenStudy (anonymous):

49 now

OpenStudy (amistre64):

i was a decent carpenter at 27 :) now my bodys broke so im going to colleget o get a masters in math :)

OpenStudy (anonymous):

been trying to figure out the Portsmouth Yardstick # for racing a non standard class sailboat in handi-cap racing

OpenStudy (anonymous):

they have formulas but not much detail on how to use them. I was a Mech Design Major so I only use math whem I have to..LOL and my old collge Calc died 2 yrs ago...guess I should go buy a new one.

OpenStudy (amistre64):

or make a new one :) it wasnt one of those wristwatch calculators was it :)

OpenStudy (anonymous):

lol no...it was a K-mart special...but a decent brand name..don't remember exactly..something like texas instruments or equiv. Couldn't afford an HP @ $300 while working way thru college. but it worked good enough to get me by with decent grades

OpenStudy (anonymous):

It even had most of the mid to high level Statistics functions so it saved my butt in Statistics..it came with a little instruction booklet on how to use it and it was very detailed on how the formulas worked...like a little cheat sheet....got A's in Stats cause of it.

OpenStudy (anonymous):

You'll have fun in Statistics..not as boring or hard as some make it look. Just remember...no matter how many times you flip a coin and it comes up heads, the odds it will be heads on the next flip is stil 50%....Statistics Trick ??? some instructors like to use.

OpenStudy (amistre64):

:) thnx and good luck